# MA-DV2F: A Multi-Agent Navigation Framework using Dynamic Velocity Vector Field

### v1

@@ III-C Estimation of Reference Speed
The reference speed $v^{(i)}_{t+1}$ is chosen after determination of the reference orientation, which depends on the situation the vehicle is in. Figure 3 shows a scenario wherein a vehicle is at the same location next to an obstacle but with opposite orientations. This determines if the velocity should move the car forward or backward.
@@
Figure 3: Shows two different scenarios when the black ego vehicle is in a collision avoiding region. In Subfigure (a), the ego vehicle is facing the obstacle, i.e. $\mathbf{u}^{T}\cdot\mathbf{X}_{obs}>0$, in which case the vehicle should be forbidden to move forwards. In Subfigure (b), the ego vehicle is oriented away from the obstacle, i.e. $\mathbf{u}^{T}\cdot\mathbf{X}_{obs}<0$, in which case the vehicle should be prevented from moving forward.
We use the logical or ($\vee$) and logical and ($\wedge$) operators [29] to describe the criteria for collision avoidance between ego vehicle $i$ and static obstacle $j$ under such situations as:
$\begin{split}F^{(i)}_{obs_{j}}&=(\alpha_{obs_{j}}^{(i)}+\epsilon_{c}\leq 0)\wedge(\gamma_{obs_{j}}^{(i)}>0)\\ B^{(i)}_{obs_{j}}&=(\alpha_{obs_{j}}^{(i)}+\epsilon_{c}\leq 0)\wedge(\gamma_{obs_{j}}^{(i)}<0)\\ \gamma^{(i)}_{obs_{j}}&={\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{X}^{(i)}_{obs_{j}}\\ \end{split}$ (6)
where $\alpha_{obs_{j}}^{(i)}$ is same as defined in Equation 3, $\epsilon_{c}$ is the an additional tolerance for the collision checking region. $F^{(i)}_{obs_{j}}$ equalling $true$ indicates the ego vehicle $i$ is forbidden to drive forwards towards obstacle $j$. This happens when the angle between the reference orientation ($\mathbf{{u}^{(i)}_{t+1}}$) and the vector from the ego vehicle to the obstacle ($\mathbf{X}^{(i)}_{obs_{j}}$) is less than $90^{\circ}$ i.e. $\gamma^{(i)}_{obs_{j}}={\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{X}^{(i)}_{obs_{j}}>0$. Likewise, $B^{(i)}_{obs_{j}}$ equaling $true$ happens when $\gamma^{(i)}_{obs_{j}}={\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{X}^{(i)}_{obs_{j}}<0$ indicating that the ego vehicle $i$ is forbidden to drive backwards.
Similarly for the case of ego vehicle $i$ and another vehicle $j$, the conditions for preventing the ego-vehicle from moving forward $F^{(i)}_{veh_{j}}$ or backward $B^{(i)}_{veh_{j}}$ are the same as described in Equation 6 except that $\mathbf{X}^{(i)}_{veh_{j}}$ replaces $\mathbf{X}^{(i)}_{obj_{j}}$.
Summarising the results above, the magnitude and sign of the velocity depends on the combination of these boolean variables i.e. $F^{(i)}_{obs_{j}}$, $F^{(i)}_{veh_{j}},B^{(i)}_{obs_{j}}$ and $B^{(i)}_{veh_{j}}$ as follows:
$\begin{split}F^{(i)}_{coll}&=(\bigvee_{j=1}^{N_{obs}}F^{(i)}_{obs_{j}})\vee(\bigvee_{j=1,j\neq i}^{N_{veh}}F^{(i)}_{veh_{j}})\\ B^{(i)}_{coll}&=(\bigvee_{j=1}^{N_{obs}}B^{(i)}_{obs_{j}})\vee(\bigvee_{j=1,j\neq i}^{N_{veh}}B^{(i)}_{veh_{j}})\\ v^{(i)}_{coll}&=\begin{cases}v_{d}&B^{(i)}_{coll}\vee(\neg F^{(i)}_{coll})\\ -v_{d}&(\neg B^{(i)}_{coll})\vee F^{(i)}_{coll}\\ 0&B^{(i)}_{coll}\vee F^{(i)}_{coll}\\ v_{tar}^{(i)}&\text{otherwise}\\ \end{cases}\end{split}$ (7)
@@
$\begin{split}v^{(i)}_{tar}&=\begin{cases}\xi^{(i)}_{p}\cdot\lambda^{(i)}_{p}\cdot v_{d}&\|\mathbf{X}^{(i)}_{tar}\|_{2}\leq r_{p}\\ v_{d}\cdot f_{sgn}({\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{\hat{u}}^{(i)}_{t+1})&\text{otherwise}\\ \end{cases}\\ \lambda^{(i)}_{p}&=\begin{cases}\bar{\lambda}^{(i)}_{p}&(\|\mathbf{X}^{(i)}_{tar}\|_{2}<\epsilon_{p})\wedge(|\theta^{(i)}_{tar}-\theta^{(i)}_{t+1}|<\epsilon_{o})\\ \sqrt{\bar{\lambda}^{(i)}_{p}}&\text{otherwise}\\ \end{cases}\\ \bar{\lambda}^{(i)}_{p}&={minimum}(\frac{\|\mathbf{X}^{(i)}_{tar}\|_{2}}{r_{p}}+\frac{|\theta^{(i)}_{tar}-\theta^{(i)}_{t+1}|}{v_{d}},1)\\ \xi^{(i)}_{p}&=\begin{cases}1&{\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{X}^{(i)}_{tar}>0.25\\ -1&{\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{X}^{(i)}_{tar}<-0.25\\ f_{sgn}(u^{(i)}_{t})&\text{otherwise}\\ \end{cases}\\ \end{split}$ (8)
@@
where $\epsilon_{p}$ and $\epsilon_{o}$ are the acceptable position and orientation tolerances for deciding to stop the vehicle at the target. When the ego vehicle is within the parking radius, it reduces its speed as it gets closer to the target position and orientation. This is achieved by the multiplier $\lambda^{(i)}_{p}$, which is proportional to $\bar{\lambda}^{(i)}_{p}$ when the vehicle state is very close to the target state and $\sqrt{\bar{\lambda}^{(i)}_{p}}$ when farther away from the tolerance. The square root accelerates the vehicle approaching its target when there is still some distance between the current state and target state, i.e. $\neg((\|\mathbf{X}^{(i)}_{tar}\|_{2}<\epsilon_{p})\wedge(|\theta^{(i)}_{tar}-\theta^{(i)}_{t+1}|<\epsilon_{o}))$. When the vehicle is already very close to the target states, i.e. $\|\mathbf{X}^{(i)}_{tar}\|_{2}<\epsilon_{p})\wedge(|\theta^{(i)}_{tar}-\theta^{(i)}_{t+1}|<\epsilon_{o}$, the ratio $\bar{\lambda}^{(i)}_{p}$ prevents the vehicle from shaking forward and backward. For $\bar{\lambda}^{(i)}_{p}$, the first term, i.e. $\frac{1}{r_{p}}\cdot\|\mathbf{X}^{(i)}_{tar}\|_{2}$, decreases linearly with the ego vehicle distance to its target, and the second term, i.e. $\frac{1}{v_{d}}\cdot|\theta^{(i)}_{tar}-\theta^{(i)}_{t+1}|$, reduces linearly with the angle difference between the reference $\theta^{(i)}_{t+1}$ and target $\theta^{(i)}_{tar}$ orientation angles. However, we do not allow the reference parking speed $v_{tar}^{(i)}$ to be any higher than the default reference speed $v_{d}$. So, $\bar{\lambda}^{(i)}_{p}$ is clipped to a maximum of $1$. $\xi^{(i)}_{p}$ controls whether the ego vehicle moves forwards or backwards by checking ${\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{X}^{(i)}_{tar}$. Originally, the vehicle should move forwards when facing the target, i.e. ${\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{X}^{(i)}_{tar}>0$, and move backwards when backing toward the target, i.e. ${\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{X}^{(i)}_{tar}<0$. However, to prevent the vehicle from changing direction at high frequency within short traveling distances, we set a margin allowing the vehicle to keep its previous direction when $|{\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{X}^{(i)}_{tar}|\leq 0.25$.
@@
These reference control commands can either be directly used to control the vehicles, or as labels to for training the GNN model in a self-supervised manner using the architecture of [12].
@@ IV-A Algorithm Evaluation
To assess the performance of our MA-DV2F approach and its learning based counterpart (self-supervised GNN model), we compare with other recent learning and search/optimization based algorithms in challenging, collision prone test cases. The recent learning based approaches include [12] which is a GNN model trained in a supervised manner and an extension of [7] i.e. GCBF+ [8], an RL-inspired GNN model incorporating safety constraints in its reward function. Meanwhile, the two search/optimization based algorithms used for comparison include CL-MAPF [9] and CSDO [10]. The test dataset comprises of challenging, collision prone scenarios with 10 vehicles - 0 obstacles to 50 vehicles - 25 obstacles, each with 1000 samples. Note that, all the static obstacles in the test cases are fixed to be circles with fixed radius because of the limitation of the CL-MAPF and CSDO environment setting and because it circumscribes an obstacle of arbitrary shape, allowing for safer behaviour. For the GCBF+, we use the DubinsCar model because it has vehicle kinematics most similar to ours. As GCBF+ has a completely different working domain with different map and agent sizes, we scale our test cases when running GCBF+ under the rule that the ratio of the map size and the agent size remains equal across all the algorithms. Details regarding the generation of test samples are given in the supplementary of the project page.
Table I shows the evaluation results of the different methods across the different vehicle-obstacle combinations described earlier against the success rate metric [8]. It measures the percentage of vehicles that successfully reach their targets within a tolerance without colliding with other objects along the way. The results show that our MA-DV2F outperforms other algorithms achieving almost 100% success rate across all vehicle-obstacle combinations. Even the self-supervised GNN model performs far better than other algorithms when scaling the number of vehicles. Only the CSDO algorithm has slightly better results than our GNN model when the number of agents are low (20). However, CSDO’s performance drops drastically as the number of agents are increased under these challenging conditions. Note that the GCBF+ pipeline only checks whether the agents reach their target positions but ignores the target orientations as shown in the project page: https://yininghase.github.io/MA-DV2F/#MC which explains why it has such poor performance. Therefore, we additionally evaluate GCBF+ by ignoring the orientation and only considering the position. Results of which are shown in the second last column. Even then, the GCBF+, does not match the performance of both our MA-DV2F and its self-supervised trained GNN counterpart.
@@ IV-B Discussion
Investigating failure Causes: Note that the Success rate metric measures a model’s ability to not only reach its destination but also avoid collisions. Therefore, for challenging test cases, a navigation algorithm may fail due to two main reasons: the algorithms behaves too aggressively by driving the vehicles to their targets albeit at the cost of colliding with other agents or behaves too conservatively to ensure safety of the vehicles resulting in some vehicles getting stuck mid-way without reaching their targets. Therefore to disambiguate between the two causes, we additionally evaluate all algorithms against the Reach rate and Safe rate metrics as proposed in [8]. Reach rate measures the percentage of vehicles that reach their targets within a tolerance disregarding any collisions along the way. Meanwhile, the Safe rate calculates the percentage of vehicles that do not collide with any other agents irrespective of whether or not they reach their targets. Figure 4 shows the performance of the different methods against the Reach Rate and Safe Rate metrics. It can be seen that the supervised GNN model behaves rather aggressively reaching the target in majority of the cases albeit at the cost of multiple collisions leading to high Reach but low Safe rates. In contrast, CL-MAPF [9] takes a greedy strategy in its optimization/search pipeline. It can quickly find the path for some vehicles which are easy to plan. But then it is unable to find paths for other vehicles and gets stuck in this partial solution. This greedy strategy lead to a low reach rate, but high safe rate since no collisions happen among vehicles that do not move.
@@
Preventing Bottlenecks: A common problem with other planning algorithms is that vehicles tend to crowd within a small sub-region in the map. This leads to these algorithms either unable to find an optimal path resulting in all vehicles becoming stationary at their place (low reach rate) or finding sub-optimal paths resulting in more collisions among vehicles (low safe rate). To prevent such bottlenecks in our MA-DV2F model, Equation 3 causes all the vehicles to drive in the counter-clockwise direction when encountering other agents. This leads to a roundabout driving behavior which breaks the bottleneck and can be visualized on the project page: https://yininghase.github.io/MA-DV2F/#RE. Due to this, all the vehicles are capable of eventually reaching their targets,by making a detour around the other agents, resulting both a high Reach and Safe rate. Our trained GNN model also adapts this behaviour.
@@
Runtime Analysis: We compare the runtime of MA-DV2F with concurrent search based methods (CSDO and CL-MAPF). TABLE II shows the total runtime of all the 1000 test cases for each vehicle-obstalce combinations. All methods were evaluated on a machine with an Intel Core i7-10750H CPU and GeForce RTX 2070 GPU. As can be seen, our MA-DV2F, is orders of magnitude faster than its peers, since it has the ability to run the different scenarios within a batch in parallel. Meanwhile CSDO and CL-MAPF are search/optimization-based algorithm that need to solve each test cases one-by-one. Note that CL-MAPF would continue its search until a solution is found, which is why the maximum allowed runtime needs to be clipped, otherwise the runtime would be even larger. Note that the evaluations in the experiments were done for only upto 50 vehicles, since other algorithms are either extremely slow or have drastically reduced performance when scaling. However, our method is capable of scaling to a far greater number of vehicles than just 50 as can be observed on the project page for scenarios with 100, 125, 250 vehicle-obstacle combinations: https://yininghase.github.io/MA-DV2F/#LE.
Lastly, note that the runtime analysis is not done for the learning based methods since, it is dependent on the GPU rather than the algorithm itself. Therefore, for the same model architecture, the runtime will be the same for all learning based algorithms

### v2

@@ III-C Estimation of Reference Speed
The reference speed $v^{(i)}_{t+1}$ is chosen after determination of the reference orientation, which depends on the situation the vehicle is in. Fig. 3(b) shows a scenario wherein a vehicle is at the same location next to an obstacle but with opposite orientations. This determines if the velocity should move the car forward or backward.
@@
Figure 3: Shows two different scenarios when the black ego vehicle is in a collision avoiding region. In Fig. 3(a)
, the ego vehicle is facing the obstacle, i.e. $\mathbf{u}^{T}\cdot\mathbf{X}_{obs}>0$, in which case the vehicle should be forbidden to move forwards. In Fig. 3(b), the ego vehicle is oriented away from the obstacle, i.e. $\mathbf{u}^{T}\cdot\mathbf{X}_{obs}<0$, in which case the vehicle should be prevented from moving backward.
We use the logical or ($\vee$) and logical and ($\wedge$) operators [29] to describe the criteria for collision avoidance between ego vehicle $i$ and static obstacle $k$ under such situations as:
$\begin{split}F^{(i)}_{obs_{k}}&=(\alpha_{obs_{k}}^{(i)}+\epsilon_{c}\leq 0)\wedge(\gamma_{obs_{k}}^{(i)}>0)\\ B^{(i)}_{obs_{k}}&=(\alpha_{obs_{k}}^{(i)}+\epsilon_{c}\leq 0)\wedge(\gamma_{obs_{k}}^{(i)}<0)\\ \gamma^{(i)}_{obs_{k}}&={\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{X}^{(i)}_{obs_{k}}\\ \end{split}$ (6)
where $\alpha_{obs_{k}}^{(i)}$ is same as defined in Equation 3, $\epsilon_{c}$ is the an additional tolerance for the collision checking region. $F^{(i)}_{obs_{k}}$ equalling $true$ indicates the ego vehicle $i$ is forbidden to drive forwards towards obstacle $k$. This happens when the angle between the reference orientation ($\mathbf{{u}^{(i)}_{t+1}}$) and the vector from the ego vehicle to the obstacle ($\mathbf{X}^{(i)}_{obs_{k}}$) is less than $90^{\circ}$ i.e. $\gamma^{(i)}_{obs_{k}}={\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{X}^{(i)}_{obs_{k}}>0$. Likewise, $B^{(i)}_{obs_{k}}$ equaling $true$ happens when $\gamma^{(i)}_{obs_{k}}={\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{X}^{(i)}_{obs_{k}}<0$ indicating that the ego vehicle $i$ is forbidden to drive backwards.
Similarly for the case of ego vehicle $i$ and another vehicle $j$, the conditions for preventing the ego-vehicle from moving forward $F^{(i)}_{veh_{j}}$ or backward $B^{(i)}_{veh_{j}}$ are the same as described in Equation 6 except that $\mathbf{X}^{(i)}_{veh_{j}}$ replaces $\mathbf{X}^{(i)}_{obs_{k}}$.
Summarising the results above, the magnitude and sign of the velocity depends on the combination of these boolean variables i.e. $F^{(i)}_{obs_{k}}$, $F^{(i)}_{veh_{j}},B^{(i)}_{obs_{k}}$ and $B^{(i)}_{veh_{j}}$ as follows:
$\begin{split}F^{(i)}_{coll}&=(\bigvee_{k=1}^{N_{obs}}F^{(i)}_{obs_{k}})\vee(\bigvee_{j=1,j\neq i}^{N_{veh}}F^{(i)}_{veh_{j}})\\ B^{(i)}_{coll}&=(\bigvee_{k=1}^{N_{obs}}B^{(i)}_{obs_{k}})\vee(\bigvee_{j=1,j\neq i}^{N_{veh}}B^{(i)}_{veh_{j}})\\ v^{(i)}_{coll}&=\begin{cases}v_{d}&B^{(i)}_{coll}\vee(\neg F^{(i)}_{coll})\\ -v_{d}&(\neg B^{(i)}_{coll})\vee F^{(i)}_{coll}\\ 0&B^{(i)}_{coll}\vee F^{(i)}_{coll}\\ v_{tar}^{(i)}&\text{otherwise}\\ \end{cases}\end{split}$ (7)
@@
$\begin{split}v^{(i)}_{tar}&=\begin{cases}\xi^{(i)}_{p}\cdot\lambda^{(i)}_{p}\cdot v_{d}&\|\mathbf{X}^{(i)}_{tar}\|_{2}\leq r_{p}\\ v_{d}\cdot f_{sgn}({\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{\hat{u}}^{(i)}_{t+1})&\text{otherwise}\\ \end{cases}\\ \lambda^{(i)}_{p}&=\begin{cases}\bar{\lambda}^{(i)}_{p}&(\|\mathbf{X}^{(i)}_{tar}\|_{2}<\epsilon_{p})\wedge(|\theta^{(i)}_{tar}-\theta^{(i)}_{t+1}|<\epsilon_{o})\\ \sqrt{\bar{\lambda}^{(i)}_{p}}&\text{otherwise}\\ \end{cases}\\ \bar{\lambda}^{(i)}_{p}&={minimum}(\frac{\|\mathbf{X}^{(i)}_{tar}\|_{2}}{r_{p}}+\frac{|\theta^{(i)}_{tar}-\theta^{(i)}_{t+1}|}{v_{d}},1)\\ \xi^{(i)}_{p}&=\begin{cases}1&{\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{X}^{(i)}_{tar}>0.25\\ -1&{\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{X}^{(i)}_{tar}<-0.25\\ f_{sgn}(\mathbf{u}^{(i)}_{t})&\text{otherwise}\\ \end{cases}\\ \end{split}$ (8)
@@
where $\epsilon_{p}$ and $\epsilon_{o}$ are the acceptable position and orientation tolerances for deciding to stop the vehicle at the target. When the ego vehicle is within the parking radius, it reduces its speed as it gets closer to the target position and orientation. This is achieved by the multiplier $\lambda^{(i)}_{p}$, which is proportional to $\bar{\lambda}^{(i)}_{p}$ when the vehicle state is very close to the target state and $\sqrt{\bar{\lambda}^{(i)}_{p}}$ when farther away from the tolerance. The square root accelerates the vehicle approaching its target when there is still some distance between the current state and target state, i.e. $\neg((\|\mathbf{X}^{(i)}_{tar}\|_{2}<\epsilon_{p})\wedge(|\theta^{(i)}_{tar}-\theta^{(i)}_{t+1}|<\epsilon_{o}))$. When the vehicle is already very close to the target state, i.e. ($\|\mathbf{X}^{(i)}_{tar}\|_{2}<\epsilon_{p})\wedge(|\theta^{(i)}_{tar}-\theta^{(i)}_{t+1}|<\epsilon_{o}$), the ratio $\bar{\lambda}^{(i)}_{p}$ prevents the vehicle from shaking forward and backward. For $\bar{\lambda}^{(i)}_{p}$, the first term: $\frac{1}{r_{p}}\cdot\|\mathbf{X}^{(i)}_{tar}\|_{2}$, decreases linearly with the ego vehicle distance to its target, and the second term, i.e. $\frac{1}{v_{d}}\cdot|\theta^{(i)}_{tar}-\theta^{(i)}_{t+1}|$, reduces linearly with the angle difference between the reference $\theta^{(i)}_{t+1}$ and target $\theta^{(i)}_{tar}$ orientation angles. However, we do not allow the reference parking speed $v_{tar}^{(i)}$ to be any higher than the default reference speed $v_{d}$. So, $\bar{\lambda}^{(i)}_{p}$ is clipped to a maximum of $1$. $\xi^{(i)}_{p}$ controls whether the ego vehicle moves forwards or backwards by checking ${\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{X}^{(i)}_{tar}$. Originally, the vehicle should move forwards when facing the target, i.e. ${\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{X}^{(i)}_{tar}>0$, and move backwards when backing toward the target, i.e. ${\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{X}^{(i)}_{tar}<0$. However, to prevent the vehicle from changing direction at high frequency within short traveling distances, we set a margin allowing the vehicle to keep its previous direction when $|{\mathbf{u}^{(i)}_{t+1}}^{T}\cdot\mathbf{X}^{(i)}_{tar}|\leq 0.25$.
@@
These reference control commands can either be directly used to control the vehicles, or as labels for training the GNN model in a self-supervised manner using the architecture of [12].
@@ IV-A Algorithm Evaluation
To assess the performance of our MA-DV2F approach and its learning based counterpart (self-supervised GNN model), we compare with other recent learning and search/optimization based algorithms in challenging, collision prone test cases. The recent learning based approaches include [12] which is a GNN model trained in a supervised manner and an extension of [7] i.e. GCBF+ [8], an RL-inspired GNN model incorporating safety constraints. Meanwhile, the two search/optimization based algorithms used for comparison include CL-MAPF [9] and CSDO [10]. The test dataset comprises of challenging, collision prone scenarios with the number of vehicles ranging from 10 to 50 and static obstacles chosen between 0 and 25. Note that, all the static obstacles in the test cases are fixed to be circles with fixed radius because of the limitation of the CL-MAPF and CSDO environment setting and because it circumscribes an obstacle of arbitrary shape, allowing for safer behaviour. For the GCBF+, we use the DubinsCar model because it has vehicle kinematics most similar to ours. As GCBF+ has a completely different working domain with different map and agent sizes, we scale our test cases when running GCBF+ under the rule that the ratio of the map size and the agent size remains equal across all the algorithms. Details regarding the generation of test samples are given in the supplementary file on the project page.
Table I shows the evaluation results of the different methods across the different vehicle-obstacle combinations described earlier against the success rate metric [8]. It measures the percentage of vehicles that successfully reach their targets within a tolerance without colliding with other objects along the way. The results show that our MA-DV2F outperforms other algorithms achieving almost 100% success rate across all vehicle-obstacle combinations. Even the self-supervised GNN model performs far better than other algorithms when scaling the number of vehicles. Only the CSDO algorithm has slightly better results than our GNN model when the number of agents are low (20). However, CSDO’s performance drops drastically as the number of agents are increased under these challenging conditions. Note that the GCBF+ pipeline only checks whether the agents reach their target positions but ignores the target orientations as shown in the project page: https://yininghase.github.io/MA-DV2F/#MC which explains why it has such poor performance. Therefore, we additionally evaluate GCBF+ by ignoring the orientation and only considering the position. Results of which are shown in the second last column. Even then, the GCBF+, does not match the performance of MA-DV2F.
@@ IV-B Discussion
Investigating failure Causes: Note that the Success rate metric measures a model’s ability to not only reach its destination but also avoid collisions. Therefore, for challenging test cases, a navigation algorithm may fail due to two main reasons: the algorithms behave too aggressively by driving the vehicles to their targets albeit at the cost of colliding with other agents or behaves too conservatively to ensure safety of the vehicles resulting in some vehicles getting stuck mid-way without reaching their targets. Therefore to disambiguate between the two causes, we additionally evaluate all algorithms against the Reach rate and Safe rate metrics as proposed in [8]. Reach rate measures the percentage of vehicles that reach their targets within a tolerance disregarding any collisions along the way. Meanwhile, the Safe rate calculates the percentage of vehicles that do not collide with any other agents irrespective of whether or not they reach their targets. Fig. 4 shows the performance of the different methods against the Reach Rate and Safe Rate metrics. It can be seen that the supervised GNN model behaves rather aggressively reaching the target in majority of the cases albeit at the cost of multiple collisions leading to high Reach but low Safe rates. In contrast, CL-MAPF [9] takes a greedy strategy in its optimization/search pipeline. It can quickly find the path for some vehicles which are easy to plan. But then it is unable to find paths for other vehicles and gets stuck in this partial solution. This greedy strategy lead to a low reach rate, but high safe rate since no collisions happen among vehicles that do not move.
@@
Preventing Bottlenecks: A common problem with other planning algorithms is that vehicles tend to crowd within a small sub-region in the map. This leads to these algorithms either unable to find an optimal path resulting in all vehicles becoming stationary at their place (low reach rate) or finding sub-optimal paths resulting in more collisions among vehicles (low safe rate). To prevent such bottlenecks in our MA-DV2F model, Equation 3 causes all the vehicles to drive in the clockwise direction when encountering other agents. This leads to a roundabout driving behavior which breaks the bottleneck and can be visualized on the project page: https://yininghase.github.io/MA-DV2F/#RE. Due to this, the vehicles are capable of eventually reaching their targets,by making a detour around the other agents, resulting in both a high Reach and Safe rate. Our trained GNN model also adapts this behaviour.
@@
Runtime Analysis: We compare the runtime of MA-DV2F with concurrent search based methods (CSDO and CL-MAPF). TABLE II shows the total runtime of all the 1000 test cases for each vehicle-obstalce combinations. All methods were evaluated on a machine with an Intel Core i7-10750H CPU and GeForce RTX 2070 GPU. As can be seen, our MA-DV2F, is orders of magnitude faster than its peers, since it has the ability to run the different scenarios within a batch in parallel. Meanwhile CSDO and CL-MAPF are search/optimization-based algorithms that need to solve each test cases one-by-one. Note that CL-MAPF would continue its search until a solution is found, which is why the maximum allowed runtime needs to be clipped, otherwise the runtime would be even larger. Note that the evaluations in the experiments were done for only upto 50 vehicles, since other algorithms are either extremely slow or have drastically reduced performance when scaling. However, our method is capable of scaling to a far greater number of vehicles than just 50 as can be observed on the project page for scenarios with 100, 125, 250 vehicle-obstacle combinations: https://yininghase.github.io/MA-DV2F/#LE.
Lastly, note that the runtime analysis is not done for the learning based methods since, it is dependent on the GPU rather than the algorithm itself. Therefore, for the same model architecture, the runtime will be the same for all learning based algorithms.
Limitations: If the vehicles are densely packed or their targets are within the safety margins of one another, then due to their non-holonomic behavior there might not be enough space for them to navigate without risking a collision. In such a scenario, the vehicles will act conservatively and hesitate to proceed so as to avoid collisions leading to less vehicles reaching the target. A similar outcome is observed, if some vehicles start behaving unexpectedly, wherein the safety margin would need to be increased to mitigate the risk of collision, albeit at the expense of reaching the target. More details on sensitivity analysis of the effect of change in static component of the safety margin ($r_{c}$) and visualization of the limitations is in the supplementary file and the project page.